How did money affect the economy of ancient Greece?
1
Expert's answer
2019-02-18T14:12:07-0500

E_d^p=((Q2-Q1)/((Q2+Q1)/2))/((P2-P1)/((P2+P1)/2))=((2000-1500)/((2000+1500)/2))/((20-25)/((20+25)/2))=0.29/(-0.22)=-1.29